# Minimal Thermodynamic Cost of Computing with Circuits

### v2

@@ IV The mismatch cost complexity of a circuit family
We aim to explore the relationship between the total MMC of a circuit and other complexity measures, such as size and depth. To this end, we extend the definition of a basis for a circuit family to include the prior distribution associated with each logic gate in the basis. For a given basis $\mathcal{B}$, we define the associated prior basis ~=$\tilde{\mathcal{B}}=\{(g,q_{g,\mathrm{pa}(g)}):g\in\mathcal{B}\}$, where $q_{g,\mathrm{pa}(g)}$ denotes the prior distribution associated with gate $g\in\mathcal{B}$. We sometimes use $q_{g}$ as a shorthand for $q_{g,\mathrm{pa}(g)}$. The definition of extended prior basis allows us to define MMC complexity associated with a circuit family.
@@
Given a basis $\mathcal{B}$ and an associated prior basis $\tilde{\mathcal{B}}$, a circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ is said to have mismatch cost complexity $m(n)$, where $m:\mathbb{N}\to\mathbb{R}^{+}$, if the associated mismatch cost $\mathcal{MC}(C_{n})$ satisfies $\mathcal{MC}(C_{n})\leq m(n)$ for every $n\in\mathbb{N}$.
@@
The MMC complexity associated with any circuit family is influenced by several key properties of the circuit. First, it depends on the basis $\tilde{\mathcal{B}}$, which includes the prior associated with each allowed logic gate in the circuit. Second, the topology of the circuit, in combination with the input distribution, jointly determines the mutual information $\mathrm{I}_{\mu}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})$, which contributes to the MMC for each gate $\mu$ downstream (see Eq. 24). Additionally, the topology and input distribution together determine the joint distribution $p_{\mu}(\boldsymbol{x}_{\mu})$ and $p_{\mu}(\boldsymbol{x}_{\mathrm{pa}}(\mu))$ for each gate $\mu$ downstream.
@@
The MMC of running a gate $\mu$ in a circuit is upper bounded by
$\mathcal{MC}_{\mu}\leq\ln\left(\frac{1}{q_{\mu}^{\min}}\right)$ (30)
where $q_{\mu}^{\min}=\min_{x}q_{\mu}(x)$, regardless of the fan-in or fan-out of the gate, and regardless of the input distribution to the gate.
Moreover, if $q_{\mu}$ is identical for all gates in all circuits in the circuit family, and $p_{in}^{n}$ is the input distribution for the circuit with input size $n$, then the total MMC of any circuit $C_{n}$ in the circuit family is upper bounded by
[…]
The proof is provided in B.1. We refer to a basis in which all prior distributions are identical for all logic gates as a homogeneous associated prior basis. Additionally, note that the term $S(p^{n}_{in})$ in Eq. 31 is upper-bounded by $n\ln 2$.
Thm. 1 establishes a relationship between the size complexity and the MMC complexity of any circuit family constructed from gates in a homogeneous prior basis. If $s(n)$ denotes the size complexity of the circuit family, then according to Eq. (31), the MMC complexity $m(n)$ of the circuit family is given by:
$m(n)\leq s(n)\ln\left(\frac{1}{q^{\min}}\right)+n\ln 2$ (32)
Figure 7: Saturating the Upper Bound on MMC for the AND Binary Tree Circuit Family (Eq. (33)). When the input is always either all ones or all zeros, then $S(p^{n}_{in})=0$ and the theoretical upper bound is $n\ln\left(1\backslash q^{\min}\right)$. The solid blue line represents the total MMC obtained by using Eq. 26 for an input distribution consisting entirely of either all ones or all zeros, matching exactly with the orange dashed line, which corresponds to the theoretical upper bound given by Eq. (33). In contrast, when the input distribution is a uniform distribution, the actual total MMC of AND binary tree remains consistently lower than the theoretical upper bound: the solid green line depicts the total MMC for a uniform input distribution, while the dashed red line shows its respective theoretical upper bound.
This upper bound is also saturable for certain circuit families with certain input distributions. For example, consider the circuit family consisting of AND binary tree. The number of gates in a binary tree with input $n$ is $|C_{n}|=n$. Therefore, according to Eq. (31), an upper bound on the MMC is
[…]
As shown in Fig. 7, when the input is always either all ones or all zeros, the theoretical upper bound is fully attained by the actual total MMC of the circuit. However, for other input distributions, such as a uniform input distribution, the upper bound is not saturated, and the actual total MMC remains consistently lower than the theoretical upper bound.
According to Eq. 32, if the size complexity $s(n)$ of a circuit family grows as a polynomial function of the input size, then the upper bound on the MMC complexity for any homogeneous prior basis also scales polynomially with the input size. In other words, any Boolean function that can be implemented by a polynomial-size circuit can also be realized by a circuit family whose MMC complexity does not exceed polynomial growth. This observation leads to the following corollary:
@@
However, it remains an open question whether the converse holds—that is, whether any function implementable by a circuit family with a MMC that grows no faster than polynomial can also be realized by a circuit family whose size complexity is polynomially bounded, and thus whether $\mathbf{P_{/poly}}=\mathbf{MMC}(\mathrm{poly})$.
Next, we examine the scenario where the prior distribution is not homogeneous across the logic gates in the basis. For a given basis $\mathcal{B}$, we denote the prior distribution associated with each gate $g\in\mathcal{B}$ by $q_{g}$. The following result tells you how the MMC complexity various with input size for an inhomogeneous prior basis:
@@
For a given basis $\mathcal{B}$, let $\{C_{n}\}_{n\in N}$ be circuit family and let and let $\tilde{\mathcal{B}}$ the associated prior basis. Then, the MMC of circuit of input size $n$ is upper bounded by,
$\mathcal{MC}(C_{n})\leq\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}+S(p_{in})$ (35)
where #$\#_{g}(n)$ is the number of gates of type $g$ in $C_{n}$ and $K_{g}=\ln\left(1\backslash q_{g}^{\min}\right).$
@@
The proof is provided in App. B.2. Note that Thm. 2 provides a slight generalization of Thm. 1 and applies to heterogeneous prior bases. Several key aspects distinguish it from the homogeneous prior case. For instance, in the homogeneous prior setting, the upper bound on the MMC is always proportional to the size complexity $|C_{n}|$ of the circuit family for every $n$. However, with heterogeneous priors, this proportionality may no longer hold.
@@
To illustrate this, consider a circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ whose basis comprises two types of gates — say, AND and XOR — labeled as $1$ and $2$. In this example, the number of AND gates scales as #$\#_{1}(n)=\log(n)$, while the number of XOR gates scales as #$\#_{2}(n)=n^{3}$. According to Thm. 2, the upper bound on the MMC is given by:
$\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}+S(p^{n}_{in})$ =#$=\#_{1}(n)K_{1}+\#_{2}(n)K_{2}+S(p^{n}_{in})$
$=\log(n)K_{1}+n^{3}K_{2}+S(p^{n}_{in})$ (36)
@@
Now, suppose the priors associated with these two types of gates are such that $K_{1}\gg K_{2}$. Under this condition, for small $n$, the upper bound on MMC scales as $\log(n)$. However, as $n$ grows large, $n^{3}$ becomes the dominant term. This example demonstrates how, in the case of heterogeneous priors, the MMC can scale differently from the size complexity of the circuit family.
@@
Figure 8: Mismatch cost scaling for a circuit family with homogeneous vs. heterogeneous priors. The circuit family that we consider here consist of binary tree made up AND and OR gates; each layer had one OR gate and rest of the gates are AND. The solid blue curve represents MMC when the prior of OR gates differs from the uniform prior of AND gates (heterogeneous priors). MMC is not linear with circuit size (Thm. 2). In contrast, the dashed orange curve represents the MMC for the same circuit family with identical priors for XOR and AND gates (homogeneous priors), and MMC scales linearly with the circuit size.
Figure 9: Mismatch cost for binary trees whose internal nodes are all 2-input AND gates, shown for depths 1–4. The three curves correspond to distinct designer priors on the gate’s input–output triple $(x,y,g)$. The agreeing prior places 97 % probability on the output that an ideal AND would produce; the uniform prior assigns equal mass to all eight states; the disagreeing prior places 97 % probability on the output opposite to AND. This clearly shows the slope of the cost vs size line depends on the prior
Figure 10: This figure illustrates that the linear cost–size relation need not hold for circuit families in the finite input regime. The figure is made by creating circuit families by “mixing and matching” small perfect-tree circuits. The first point is a depth-1 tree. For the depth-2 tree (three internal gates) we pick the truth table with the smallest per-gate slope under a common skewed prior $Q_{par}(0,0)=0.9$,$Q_{g}(0)=0.9$. The depth-3 tree uses the second-smallest-slope gate, and the depth-4 tree uses the gate with the largest slope. Although each individual segment is linear in circuit size, the deliberate change of gate type at each depth yields a piecewise-linear curve whose overall shape is distinctly non-linear, illustrating how mixing gate families can steer the global cost profile even when the prior is held fixed for all gates
In Fig. 10, we demonstrate how, due to the effect of heterogeneous prior distributions, the mismatch cost can deviate from the linear scaling with size.
The lower bound on the mismatch cost
Going back to Eq. (24) for MMC of a gate in a circuit, under the assumption that the prior distribution $q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\mathrm{pa}(\mu))$ is a product distribution, $q_{\mu}(\boldsymbol{x}_{\mu})q_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$, the expression for MMC simplifies to:
$\mathcal{MC}_{\mu}=\mathrm{I}_{0}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{ch}(\mu)})+D\left(p_{\mu}\|q_{\mu}\right)$ (37)
It turns out that for certain kinds of circuits, the mutual information between the gate and its children is zero, e.g., in an XOR binary tree. Moreover, if the actual distribution $p_{\mu}$ perfectly matches the prior $q_{\mu}$, the KL-divergence term becomes zero. Therefore, it is theoretically possible for MMC of a gate to approach zero.
This bound can even be achieved for certain types of circuits with specific prior distributions. For example, in an XOR binary tree fed with a uniform input distribution, the distribution $p_{\mu}$ remains a uniform distribution for every downstream XOR gate in the circuit. If all the XOR gates have a uniform prior, $q_{\mu}$, then the actual distribution $p_{\mu}$ perfectly matches with the prior, leading to zero KL divergence and, hence, zero MMC.
However, even if the MMC of each gate in the circuit is zero, the total MMC includes the overwriting mismatch cost, which is lower bounded by the mutual information between input and non-input gates. Therefore, total MMC is lower bounded by the mutual information between the input and non-input gates,
$\mathcal{MC}(C_{n})\geq\mathrm{I}_{0}(\mathrm{X}_{in},\mathrm{X}_{nin})$ (38)
For a deterministic circuit, the input uniquely determines the state of non-input gates. Therefore, the mutual information simplifies to the entropy of the input distribution.
@@ V.1 MMC of different circuit families implementing the same function family
The problem of computing a Boolean function with a circuit family under resource constraints naturally leads to the challenge of optimizing both circuit size and depth complexity. Since a Boolean function can be computed by multiple circuit families, size and depth serve as key measures of computational efficiency. The goal, therefore, is to identify the most efficient circuit family—one that minimizes both size and depth complexity. While size complexity corresponds to space efficiency and depth complexity to time efficiency, mismatch cost complexity serves as a measure of thermodynamic efficiency—the minimal dissipated heat in a circuit. Thus, when comparing two circuit families computing the same Boolean function, MMC is a key metric for evaluating thermodynamic efficiency.
Figure 11: MMC of two different circuit families implementing the addition function, which takes two integers in the standard $n$-bit representation and returns their sum in the standard $(n+1)$-bit representation.
@@
For example, the Boolean function $\text{ADD}^{n}$, which computes the sum of two $n$-bit binary numbers as discussed in Sec. II.1, can be implemented by two circuit families: the ripple-carry adder and the carry look-ahead adder. RCA has linear size and depth complexity, while CLA has size complexity of $n\log n$ and logarithmic depth complexity.
@@
In Fig. 11, we compare MMC of RCA and CLA across various input sizes. The results show a linear scaling of mismatch cost with size complexity for both circuits. However, as input size increases, RCA consistently exhibits a lower MMC than CLA, indicating that despite being slower, RCA is thermodynamically more efficient.
V.2 MMC of a circuit family for heterogeneous priors

### v3

@@ IV The mismatch cost complexity of a circuit family
We aim to explore the relationship between the total mismatch cost of a circuit and other complexity measures, such as size and depth. To this end, we extend the definition of a basis for a circuit family to include the prior distribution associated with each logic gate in the basis. For a given basis $\mathcal{B}$, we define the associated prior basis $\tilde{\mathcal{B}}=\{(g,q_{g,\mathrm{pa}(g)}):g\in\mathcal{B}\}$, where $q_{g,\mathrm{pa}(g)}$ denotes the prior distribution associated with gate $g\in\mathcal{B}$. The definition of extended prior basis allows us to define mismatch cost complexity associated with a circuit family.
@@
Given a basis $\mathcal{B}$ and an associated prior basis $\tilde{\mathcal{B}}$, a circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ is said to have mismatch cost complexity $m(n)$, where $m:~\mathbb{N}~\to~\mathbb{R}^{+}$, if its mismatch cost is uniformly bounded in the worst case over input distributions, i.e.,
$\mathcal{MC}(C_{n},p_{\mathrm{in}})\leq m(n)$ (31)
for all input distributions $p_{\mathrm{in}}$ and all $n\in\mathbb{N}$.
@@
The mismatch cost complexity associated with any circuit family is influenced by several key properties of the circuit. For example, it depends on the basis $\tilde{\mathcal{B}}$, which includes the types of logic gates allowed in the circuit and their associated prior distribution. Moreover, the topology of the circuit, in combination with the input distribution, jointly determines the mutual information $\mathrm{I}_{\mu+1}(\mathrm{X}_{\mu};\mathrm{X}_{\mathrm{pa}(\mu)})$, which contributes to the mismatch cost for each gate $\mu$ (see Eq. (25)). Additionally, the topology and input distribution together determine the distributions $p_{\mu}(\boldsymbol{x}_{\mu})$ and $p_{\mu}(\boldsymbol{x}_{\mathrm{pa}}(\mu))$ for each gate $\mu$.
@@
Suppose that:
  1. (a)
    For every gate $\mu$ in the circuit, the prior distribution over $\mu$ and its parents $\mathrm{pa}(\mu)$ factorizes as a product distribution; $q_{\mu,\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mu},\boldsymbol{x}_{\mathrm{pa}(\mu)})=q_{\mu}(\boldsymbol{x}_{\mu})q_{\mathrm{pa}(\mu)}(\boldsymbol{x}_{\mathrm{pa}(\mu)})$
  2. (b)
    All gates share the same marginal prior, i.e., $q_{\mu}=q_{g}$ for all gate types $g$,
  3. (c)
    The prior distribution over the $n$-input nodes $(x_{\mathrm{in},1}\ldots x_{\mathrm{in},n})$ is also a product distribution; $q_{\mathrm{in}}(\boldsymbol{x}_{\mathrm{in}})=\prod_{k=1}^{n}q_{i}(x_{\mathrm{in},k})$
Then the total mismatch cost of running the circuit $C_{n}$, in the worst case over input distributions, satisfies the upper bound
$\mathcal{MC}(C_{n})\leq|C_{n}|\,K+n\,K_{\mathrm{in}},$ (32)
independently of the fan-in and fan-out of the gates, where
$K=\log\!\left(\frac{1}{q_{g}^{\min}}\right),\qquad K_{\mathrm{in}}=\log\!\left(\frac{1}{q_{i}^{\min}}\right)$
are positive constants determined by the minimal components of the gate and input node priors, respectively.
The proof of Theorem 1 is provided in (B.1). We refer to a basis in which all the logic gates have identical prior distributions as a homogeneous associated prior basis. Thm. (1) establishes a relationship between the size complexity and the mismatch cost complexity of any circuit family constructed from gates in a homogeneous prior basis. If $s(n)$ denotes the size complexity of the circuit family, then according to Eq. (32), the mismatch cost complexity $m(n)$ of the circuit family is given by:
$m(n)\leq s(n)\,K+n\,K_{\mathrm{in}},$ (33)
Figure 7: Saturating the Upper Bound on MMC for the AND Binary Tree Circuit Family. The solid blue line represents the total mismatch cost obtained by using Eq. (25) and a prior that is uniform for an input distribution consisting entirely of all ones, matching exactly with the orange dashed line, which corresponds to the theoretical upper bound given by Eq. (32). In contrast, when the input distribution is a uniform distribution, the actual total mismatch cost of AND binary tree remains consistently lower than the theoretical upper bound: the solid green line depicts the total mismatch cost for a uniform input distribution.
According to Eq. (33), if the size complexity $s(n)$ of a circuit family grows as a polynomial function of the input size, then the upper bound on the mismatch cost complexity for any homogeneous prior basis also scales polynomially with the input size. In other words, any Boolean function that can be implemented by a polynomial-size circuit can also be realized by a circuit family whose mismatch cost complexity does not exceed polynomial growth. This observation leads to the following corollary:
@@
It remains an open question whether the converse holds—that is, whether any function implementable by a circuit family with a mismatch cost that grows no faster than polynomial can also be realized by a circuit family whose size complexity is polynomial bounded, and thus whether $\mathbf{P_{/poly}}=\mathbf{MMC}(\mathrm{poly})$.
The worst-case upper bound in Theorem 1 can be saturated for certain circuit families with certain (worst-case) input distributions. For example, consider the circuit family consisting of AND binary tree. The number of gates in a binary tree with input $n$ is $|C_{n}|=n$. Therefore, according to Eq. (32), an upper bound on the MMC is
$\mathcal{MC}(C_{n})\leq n\log\left(\frac{1}{q^{\min}}\right)+n\log\left(\frac{1}{q_{i}^{\min}}\right)$ (35)
As shown in Fig. 7, when the input is always all ones, the theoretical upper bound is fully attained by the actual total MMC of the circuit. However, for other input distributions, such as a uniform input distribution, the upper bound is not saturated, and the actual total mismatch cost remains consistently lower than the theoretical upper bound.
Next, we examine the scenario where the prior distribution is not homogeneous across the logic gates in the basis, i.e., when assumption (b) of Thm. (1) no longer holds. For a given basis $\mathcal{B}$, we denote the prior distribution associated with each gate $g\in\mathcal{B}$ by $q_{g}$. The following result tells you how the mismatch cost complexity varies with input size for an inhomogeneous prior basis:
@@
For a given basis $\mathcal{B}$, let $\{C_{n}\}_{n\in N}$ be circuit family and let $\tilde{\mathcal{B}}=\{(g,q_{g})|g\in\mathcal{B}\}$ be the associated prior basis. Then, the mismatch cost of circuit of input size $n$ is upper bounded by,
$\mathcal{MC}(C_{n},p_{\mathrm{\mathrm{in}}})\leq\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}+nK_{\mathrm{in}}.$ (36)
where $_{g}(n)$ is the number of gates of type $g$ in $C_{n}$ and $K_{g}=\log\left(1\backslash q_{g}^{\min}\right).$
@@
The proof is provided in App. B.2. Note that Thm. (2) provides a slight generalization of Thm. (1) and applies when different logic gates have different priors. Several key aspects distinguish it from the homogeneous prior case. For instance, in the homogeneous prior setting, the worst-case upper bound on the mismatch cost is always proportional to the size complexity $|C_{n}|$ of the circuit family for every $n$. However, with heterogeneous priors, this proportionality may no longer hold.
@@
To illustrate this, consider a circuit family $\{C_{n}\}_{n\in\mathbb{N}}$ whose basis comprises two types of gates — say, AND and XOR — labeled as $1$ and $2$. Assume that the number of AND gates in this circuit family scales as $_{1}(n)=\log(n)$, while the number of XOR gates scales as $_{2}(n)=n^{3}$. According to Thm. (2), the upper bound on the mismatch cost is given by:
$\sum_{g\in\mathcal{B}}\#_{g}(n)K_{g}+nK_{\mathrm{in}}$ $=\#_{1}(n)K_{1}+\#_{2}(n)K_{2}+nK_{\mathrm{in}}$
$=\log(n)K_{1}+n^{3}K_{2}+nK_{\mathrm{in}}$ (37)
@@
Now, suppose the priors associated with these two types of gates are such that $K_{1}\gg K_{2}$ and $K_{1}\gg K_{\mathrm{in}}$. This can arise, for example, when $q_{1}^{\min}<q_{2}^{\min}$. Under this condition, for small $n$, the upper bound on the mismatch cost scales as $\log(n)$. However, as $n$ grows large, $n^{3}$ becomes the dominant term. This example demonstrates how, in the case of heterogeneous priors, the MMC can scale differently from the size complexity of the circuit family.
Figure 8: The MMC of two different circuit families implementing the addition function, which takes two integers in the standard $n$-bit representation and returns their sum in the standard $(n+1)$-bit representation.
@@
Figure 9: Mismatch cost scaling for a circuit family with homogeneous (orange) vs. heterogeneous priors (blue). The circuit family that we consider here consist of binary tree made up of AND and OR gates; each layer had one OR gate and rest of the gates are AND. The solid blue curve shows the mismatch cost when the prior of OR gates differs from the uniform prior of AND gates (heterogeneous priors). For the case where AND and OR gates have different prior distribution, the mismatch cost is not linear with circuit size (Thm. (2)). In contrast, the orange line represents the mismatch cost for the same circuit family with identical priors for OR and AND gates (homogeneous priors), and the mismatch cost scales linearly with the circuit size.
Figure 10: Mismatch cost for a family of binary-tree AND circuits under different gate priors. For a uniform input distribution, the evaluation of the circuit induces a joint distribution over output $\boldsymbol{x}_{g}$ and inputs $\boldsymbol{x}_{\mathrm{pa}(g)}$ of each gate $g$. The figure shows the total mismatch cost for three choices of the gate prior $q_{g}(\boldsymbol{x}_{g},\boldsymbol{x}_{\mathrm{pa}(g)})$: (blue) a prior ‘agreeing’ to the ideal AND behavior, assigning 97% probability to the correct AND output given the inputs; (orange) a uniform prior over the joint input–output states; (green) a prior that assigns 97% probability to the output opposite to that of an ideal AND gate (‘disagreeing’ with the ideal AND behavior). The comparison illustrates how agreement or disagreement between the prior and the actual gate behavior impacts the total mismatch cost.
Figure 11: Mismatch cost of a mixed binary tree circuit family. This figure illustrates that the linear cost–size relation need not hold for circuit families in the finite input regime. The circuit family is constructed as follows: starting with depth-1 tree which consists of an AND gate, the depth-2 tree is made of CONST gates that always output zero, depth-3 tree made from NAND gates and lastly, depth 4 tree is a tree of OR gates. All gates are evaluated under the same prior: inputs are assumed independent and uniformly distributed, so that $q_{\mathrm{pa}(g)}(\boldsymbol{x}_{\mathrm{pa}}(g))=\tfrac{1}{4}$ for all $\boldsymbol{x}_{\mathrm{pa}}(g)\in\{0,1\}^{2}$, and $q_{g|\mathrm{pa}(g)}(\boldsymbol{x}_{g}\mid\boldsymbol{x}_{\mathrm{pa}}(g))~=~\begin{cases}1-\varepsilon&\text{if }g=x\land y,\\ \varepsilon&\text{otherwise},\end{cases}$, with $\varepsilon=10^{-12}$. This means that the prior strongly favors AND behavior, and deviations from AND incur mismatch cost. Jointly, $q_{\mathrm{pa}(g)}$ and $q_{g|\mathrm{pa}(g)}$ determine the prior $q_{g,~\mathrm{pa}(g)}(\boldsymbol{x}_{g},~\boldsymbol{x}_{\mathrm{pa}(g)})~=~q_{g|\mathrm{pa}(g)}(\boldsymbol{x}_{g}~\mid~\boldsymbol{x}_{\mathrm{pa}}(g))q_{\mathrm{pa}(g)}(\boldsymbol{x}_{\mathrm{pa}}(g))$.
In Fig. 9, we illustrate how heterogeneous prior distributions across gate types can lead to deviations from linear scaling of mismatch cost with circuit size. Specifically, Fig. 9 considers a family of binary-tree circuits composed of AND and OR gates, where each layer contains a single OR gate and the remaining gates are AND gates. The prior distributions for the AND gates are fixed to be uniform, while the prior assigned to the OR gates is allowed to vary. The figure compares two scenarios: one in which the OR gates share the same prior as the AND gates, and another in which the OR gates have a different prior. In the former case, the total MMC scales linearly with circuit size, whereas in the latter case this linear scaling is broken. This example illustrates how heterogeneity in gate priors alone—without any change in circuit topology—can induce deviations from linear MMC scaling.
@@ V.1 MMC of different circuit families implementing the same function family
The problem of computing a Boolean function with a circuit family under resource constraints naturally leads to the challenge of optimizing both circuit size and depth complexity. Since a Boolean function can be computed by multiple circuit families, size and depth serve as key measures of computational efficiency. The goal, therefore, is to identify the most efficient circuit family—one that minimizes both size and depth complexity. While size complexity corresponds to space efficiency and depth complexity to time efficiency, mismatch cost complexity serves as a measure of thermodynamic efficiency—the minimal dissipated heat in a circuit. Thus, when comparing two circuit families computing the same Boolean function, mismatch cost is a key metric for evaluating thermodynamic efficiency.
@@
For example, the Boolean function $\text{ADD}^{n}$, which computes the sum of two $n$-bit binary numbers (discussed in Sec. II.1), can be implemented by two circuit families: the ripple-carry adder and the carry look-ahead adder. RCA has linear size and depth complexity, while CLA has size complexity of $n\log n$ and logarithmic depth complexity.
@@
In Fig. 8, we compare the mismatch cost of RCA and CLA across various input sizes. The results show a linear scaling of mismatch cost with size complexity for both circuits. However, as input size increases, RCA consistently exhibits a lower mismatch cost than CLA, indicating that despite being slower, RCA is thermodynamically more efficient.
V.2 Relationship between MMC complexity and size complexity
Theorems 1 and 2 provide an upper bound on the total mismatch cost (MMC) as a function of the number of gates in a circuit. This bound is worst-case over all input distributions, meaning it holds regardless of the specific inputs fed to the circuit.
The relationship between MMC and circuit size can be further explored under different choices of prior distributions. In the case of a homogeneous prior basis, all logical gates in the circuit share the same prior distribution. The particular choice of this prior determines the slope of the resulting linear relationship between MMC and circuit size. For instance, Fig. 10 shows the MMC for a fixed circuit family of AND binary trees under several prior distributions; as expected, the slope of the linear relationship changes with the prior.
It is important to note that the bounds in Theorem 1 are worst-case. In practice, the actual MMC may not scale linearly with circuit size for certain circuit families under specific input distributions. As an example, Fig. 11 illustrates a circuit family for which the MMC exhibits a highly non-linear dependence on circuit size.